# An Adaptive Approach for Infinitely Many-armed Bandits under Generalized Rotting Constraints

### v3

@@ 3 Algorithms and Regret Analysis
Here we consider the case of slow rotting, where, recall, the adaptive adversary is constrained such that the total amount of rotting is bounded by $V_{T}$. We analyze the regret of Algorithm 1 with tuned $\delta$ using $\beta$ and $V_{T}$. We define $\delta_{V}(\beta)=\max\{(V_{T}/T)^{1/(\beta+2)},1/T^{1/(\beta+1)}\}$ when $\beta\geq 1$ and $\delta_{V}(\beta)=\max\{(V_{T}/T)^{1/3},1/\sqrt{T}\}$ when $10<\beta<$ 1. The algorithm with $\delta_{V}(\beta)$ achieves a regret bound in the following theorem.
@@
Next, we provide a bound for $\mathbb{E}[R^{\mathcal{B}}(T)]$. We employ episodic regret analysis, defining an episode as the time steps between consecutively selected distinct good arms by the algorithm. By analyzing bad arms within each episode, we can derive an upper bound for the overall regret arising from bad arms. We define the regret from bad arms over $m^{\mathcal{G}}$ episodes as $R^{\mathcal{B}}_{m^{\mathcal{G}}}$. We first consider the case of $V_{T}=\omega(\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\})$. In this case, by setting $m^{\mathcal{G}}=\lceil 2V_{T}/\delta\rceil$, we can show that $R^{\mathcal{B}}(T)\leq R_{m^{\mathcal{G}}}^{\mathcal{B}}$ with a high probability. By analyzing $R_{m^{\mathcal{G}}}^{\mathcal{B}}$ with the episodic analysis, we can show that $\mathbb{E}[R^{\mathcal{B}}(T)]\leq\mathbb{E}[R^{\mathcal{B}}_{m^{\mathcal{G}}}]=\tilde{O}(\max\{T^{\frac{\beta+1}{\beta+2}}V_{T}^{\frac{1}{\beta+2}},T^{\frac{2}{3}}V_{T}^{\frac{1}{3}}\}).$ As in the similar manner, when $V_{T}=O(\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\})$, by setting $m^{\mathcal{G}}=C_{3}$ for some constant $C_{3}>0$, we can show that $\mathbb{E}[R^{\mathcal{B}}(T)]\leq\mathbb{E}[R^{\mathcal{B}}_{m^{\mathcal{G}}}]=\tilde{O}(\max\{T^{\frac{\beta}{\beta+1}},\sqrt{T}\}).$ From the above two inequalities, we have
@@
Here we consider abruptly rotting reward distribution under the constraint of $S_{T}$. We consider Algorithm 1 with $\delta$ newly tuned by $S_{T}$ and $\beta$. We define $\delta_{S}(\beta)=(S_{T}/T)^{1/(\beta+1)}$ for $\beta\geq 1$ and $\delta_{S}(\beta)=(S_{T}/T)^{1/2}$ for $0<\beta\leq 1$. We also define $\bar{V}_{T}=\sum_{t=1}^{T-1}\mathbb{E}[\rho_{t}]$. In the following theorem, we present a regret upper bound for Algorithm 1 with $\delta_{S}(\beta)$.
@@
Here we study the case when the algorithm does not have prior information about the values of $V_{T}$, $S_{T}$, and $\beta$ under the constraints of $V_{T}$ and $S_{T}$. These parameters play a crucial role in determining the optimal threshold parameter $\delta$ in Algorithm 1. We propose an algorithm based on estimating the optimal threshold parameter $\delta$ directly (Algorithm 2), rather than estimating each unknown parameter separately, employing the Bandit-over-Bandit (BoB) approach [11]. Under assumptions concerning the bounds for the cumulative amount of rotting and a constrained version of the adaptive adversary for rotting rates, which are less general than Assumptions 2.1 and 2.4 but still more general than those in Kim et al. [13], the algorithm achieves a regret bound of $\mathbb{E}[R^{\pi}(T)]=\tilde{O}(\min\left\{R_{V},R_{S}\right\}+\max\{T^{(2\beta+1)/(2\beta+2)},\sqrt{T}\})$. The additional cost arises from learning $\delta$ compared to the regret bound of Corollary 3.4. Further details of the algorithm and regret analysis are provided in Appendix A.6.
@@ 5 Experiments
We first compare the performance of our Algorithms 1 and 2 with UCB-TP [13], the state-of-the-art algorithm for the rotting setting, and SSUCB [5], a near-optimal algorithm for stationary infinitely many-armed bandits. For comparison with UCB-TP, recall our discussion in Remark 3.2. We set the rotting rates such that $\rho_{t}=1/(t\log(T))$ for all $t$, for which $\rho=\rho_{1}=1/\log(T)=o(1)$, $V_{T}=O(1)$, and $S_{T}=T$. In Figure 5 (a), we can observe that Algorithms 1 and 2 perform better than UCB-TP and SSUCB (and Algorithm 1 outperforms Algorithm 2), which is in agreement with our theoretical analysis for the case $\beta=1$. In this case, the regret bounds for Algorithms 1 and 2 are $\tilde{O}(T^{2/3})$ and $\tilde{O}(T^{3/4})$ from Corollary 3.4 and Theorem A.15, respectively, which are tighter than the regret bound of $\tilde{O}(T/\log(T)^{1/3})$ for UCB-TP.
We next compare the performance of our Algorithms with benchmarks for smaller or larger $\beta$. In Figure 5 (b,c), we can observe that our algorithms outperform the benchmarks for $\beta=0.5$ and $\beta=2$. Additional experiments can be found in Appendix A.10.
@@ 7 Acknowledgements
JK was supported by the Global-LAMP Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2023-00301976). SY was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2022-II220311, Development of Goal-Oriented Reinforcement Learning Techniques for Contact-Rich Robotic Manipulation of Everyday Objects)
@@ A.1 Limitations & Discussion
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $2\delta$, given a fixed value of $m^{\mathcal{G}}$. Therefore, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-2\delta)^{k}2\delta$, for $i=1,\ldots,m^{\mathcal{G}}$. Define $\tilde{m}_{T}$ to be the number of episodes from the policy $\pi$ over the horizon $T$, $\tilde{m}_{T}^{\mathcal{G}}$ to be the total number of samples of a good arm by the policy $\pi$ over the horizon $T$ such that $\tilde{m}_{T}^{\mathcal{G}}=\tilde{m}_{T}$ or $\tilde{m}_{T}^{\mathcal{G}}=\tilde{m}-1$, and $\tilde{m}_{i,T}^{\mathcal{B}}$ to be the number of samples of a bad arm in the $i$-th episode by the policy $\pi$ over the horizon $T$.
@@
In this case, we have $\delta=\delta_{V}(\beta)=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}.$ Here, we define the policy $\pi$ after time $T$ such that it pulls a good arm until its total rotting variation is equal to or greater than $\delta/2$ and does not pull a sampled bad arm. We note that defining how $\pi$ works after $T$ is only for the proof to get a regret bound over time horizon $T$. For the last arm $\tilde{a}$ over the horizon $T$, it pulls the arm until its total variation becomes $\max\{\delta/2,V_{T}(\tilde{a})\}$ if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the total rotting variation of pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
$=\mathbb{P}(\Delta_{1}(a)=x)/(1-C(2\delta)^{\beta}).$
We note that $2\delta<\Delta_{t_{1}}(a(i,j))=\Delta_{1}(a(i,j))\leq 1$. Since $n_{i,j}^{\mathcal{B}}=\tilde{O}(1/(\Delta_{t_{1}}(a(i,j))-\delta)^{2})=\tilde{O}(1/\delta^{2})$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ and $m^{\mathcal{G}}=\lceil 2V_{T}/\delta\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta)^{\beta}-1$, we have
@@
Case 2: Now we consider $V_{T}\leq\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\}$ in the following. In this case, we have $\delta=\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}.$ For getting $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}$, here we define the policy $\pi$ after time $T$ such that it pulls $V_{T}$ amount of rotting variation for a good arm and 00 for a bad arm. We note that defining how $\pi$ works after $T$ is only for the proof to get a regret bound over time horizon $T$. For the last arm $\tilde{a}$ over the horizon $T$, it pulls the arm up to $V_{T}$ amount of rotting variation if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the amount of rotting variation from pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
Under $E_{1}$, when $m^{\mathcal{G}}=3$, we have R^B(T)R^B_m^G.
@@
From Lemma A.2, under $E_{1}$ we can find that $V_{i}^{\mathcal{G}}\geq\min\{\delta/2,V_{T}\}$ for $i\in[m^{\mathcal{G}}]$. Then if $m^{\mathcal{G}}=3$, then with $\delta=\max\{1/T^{1/(\beta+1)},1/\sqrt{T}\}$ and $V_{T}\leq\max\{1/T^{1/(\beta+1)},1/\sqrt{T}\}$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=\max\{(1/T)^{1/(\beta+1)},1/T^{1/2}\}$ and $m^{\mathcal{G}}=C_{3}$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta)^{\beta}-1$, we have
@@
From Lemma A.1, with $\delta=\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}$ we have
@@
From $R^{\pi}(T)=R^{\mathcal{G}}(T)+R^{\mathcal{B}}(T)$ and Lemmas A.1, A.5, A.6 with $\delta=\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}$ we have

### v4

@@ 3 Algorithms and Regret Analysis
Here we consider the case of slow rotting, where, recall, the adaptive adversary is constrained such that the total amount of rotting is bounded by $V_{T}$. We analyze the regret of Algorithm 1 with tuned $\delta$ using $\beta$ and $V_{T}$. We define $\delta_{V}(\beta)=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},1/T^{1/(\beta+1)}\}$ when $\beta\geq 1$ and $\delta_{V}(\beta)=c_{1}\max\{(V_{T}/T)^{1/3},1/\sqrt{T}\}$ when $10<\beta<$ 1 for some constant $10<c_{1}<$ 1. The algorithm with $\delta_{V}(\beta)$ achieves a regret bound in the following theorem.
@@
Next, we provide a bound for $\mathbb{E}[R^{\mathcal{B}}(T)]$. We employ episodic regret analysis, defining an episode as the time steps between consecutively selected distinct good arms by the algorithm. By analyzing bad arms within each episode, we can derive an upper bound for the overall regret arising from bad arms. We define the regret from bad arms over $m^{\mathcal{G}}$ episodes as $R^{\mathcal{B}}_{m^{\mathcal{G}}}$. We first consider the case of $V_{T}>\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\}$. In this case, by setting $m^{\mathcal{G}}=\lceil 2V_{T}/\delta\rceil$, we can show that $R^{\mathcal{B}}(T)\leq R_{m^{\mathcal{G}}}^{\mathcal{B}}$ with a high probability. By analyzing $R_{m^{\mathcal{G}}}^{\mathcal{B}}$ with the episodic analysis, we can show that $\mathbb{E}[R^{\mathcal{B}}(T)]\leq\mathbb{E}[R^{\mathcal{B}}_{m^{\mathcal{G}}}]=\tilde{O}(\max\{T^{\frac{\beta+1}{\beta+2}}V_{T}^{\frac{1}{\beta+2}},T^{\frac{2}{3}}V_{T}^{\frac{1}{3}}\}).$ As in the similar manner, when $V_{T}\leq\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\}$, by setting $m^{\mathcal{G}}=C_{3}$ for some constant $C_{3}>0$, we can show that $\mathbb{E}[R^{\mathcal{B}}(T)]\leq\mathbb{E}[R^{\mathcal{B}}_{m^{\mathcal{G}}}]=\tilde{O}(\max\{T^{\frac{\beta}{\beta+1}},\sqrt{T}\}).$ From the above two inequalities, we have
@@
Here we consider abruptly rotting reward distribution under the constraint of $S_{T}$. We consider Algorithm 1 with $\delta$ newly tuned by $S_{T}$ and $\beta$. We define $\delta_{S}(\beta)=c_{1}(S_{T}/T)^{1/(\beta+1)}$ when $\beta\geq 1$ and $\delta_{S}(\beta)=c_{1}(S_{T}/T)^{1/2}$ when $0<\beta\leq 1$ for some constant $10<c_{1}<$ 1. We also define $\bar{V}_{T}=\sum_{t=1}^{T-1}\mathbb{E}[\rho_{t}]$. In the following theorem, we present a regret upper bound for Algorithm 1 with $\delta_{S}(\beta)$.
@@
Here we study the case when the algorithm does not have prior information about the values of $V_{T}$, $S_{T}$, and $\beta$ under the constraints of $V_{T}$ and $S_{T}$. These parameters play a crucial role in determining the optimal threshold parameter $\delta$ in Algorithm 1. We propose an algorithm based on estimating the optimal threshold parameter $\delta$ directly (Algorithm 2), rather than estimating each unknown parameter separately, employing the Bandit-over-Bandit (BoB) approach [11]. Under assumptions concerning the bounds for the cumulative amount of rotting and a constrained version of the adaptive adversary for rotting rates, which are less general than Assumptions 2.1 and 2.4 but still more general than those in Kim et al. [13], the algorithm achieves a regret bound of $\mathbb{E}[R^{\pi}(T)]=\tilde{O}(\min\left\{R_{V},R_{S}\right\}+\max\{T^{(2\beta+1)/(2\beta+2)},T^{3/4}\})$. The additional cost arises from learning $\delta$ compared to the regret bound of Corollary 3.4. Further details of the algorithm and regret analysis are provided in Appendix A.6.
@@ 5 Experiments
We first compare the performance of our Algorithms 1 and 2 with UCB-TP [13], the state-of-the-art algorithm for the rotting setting, and SSUCB [5], a near-optimal algorithm for stationary infinitely many-armed bandits. For comparison with UCB-TP, recall our discussion in Remark 3.2. We set the rotting rates such that $\rho_{t}=1/(t\log(T))$ for all $t$, for which $\rho=\rho_{1}=1/\log(T)=o(1)$, $V_{T}=O(1)$, and $S_{T}=T$. In Figure 3, we can observe that Algorithms 1 and 2 perform better than UCB-TP and SSUCB (and Algorithm 1 outperforms Algorithm 2), which is in agreement with our theoretical analysis for the case $\beta=1$. In this case, the regret bounds for Algorithms 1 and 2 are $\tilde{O}(T^{2/3})$ and $\tilde{O}(T^{3/4})$ from Corollary 3.4 and Theorem A.15, respectively, which are tighter than the regret bound of $\tilde{O}(T/\log(T)^{1/3})$ for UCB-TP. Additional experiments can be found in Appendix A.10.
@@ 7 Acknowledgements
The authors thank Joe Suk and the anonymous reviewers for helpful discussions. JK was supported by the Global-LAMP Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2023-00301976). SY was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2022-II220311, Development of Goal-Oriented Reinforcement Learning Techniques for Contact-Rich Robotic Manipulation of Everyday Objects)
@@ A.1 Limitations & Discussion
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $2\delta$, given a fixed value of $m^{\mathcal{G}}$. Therefore, with some constant $C>0$, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-C(2\delta)^{\beta})^{k}C(2\delta)^{\beta}$, for $i=1,\ldots,m^{\mathcal{G}}$. Define $\tilde{m}_{T}$ to be the number of episodes from the policy $\pi$ over the horizon $T$, $\tilde{m}_{T}^{\mathcal{G}}$ to be the total number of samples of a good arm by the policy $\pi$ over the horizon $T$ such that $\tilde{m}_{T}^{\mathcal{G}}=\tilde{m}_{T}$ or $\tilde{m}_{T}^{\mathcal{G}}=\tilde{m}-1$, and $\tilde{m}_{i,T}^{\mathcal{B}}$ to be the number of samples of a bad arm in the $i$-th episode by the policy $\pi$ over the horizon $T$.
@@
In this case, we have $\delta=\delta_{V}(\beta)=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}.$ Here, we define the policy $\pi$ after time $T$ such that it pulls a good arm until its total rotting variation is equal to or greater than $\delta/2$ and does not pull a sampled bad arm. We note that defining how $\pi$ works after $T$ is only for the proof to get a regret bound over time horizon $T$. For the last arm $\tilde{a}$ over the horizon $T$, it pulls the arm until its total variation becomes $\max\{\delta/2,V_{T}(\tilde{a})\}$ if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the total rotting variation of pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
$=\mathbb{P}(\Delta_{1}(a)=x)/(1-C(2\delta)^{\beta}),$
where $C(2\delta)^{\beta}<1$ with small enough positive constant $c_{1}<1$ for $\delta$. We note that $2\delta<\Delta_{t_{1}}(a(i,j))=\Delta_{1}(a(i,j))\leq 1$. Since $n_{i,j}^{\mathcal{B}}=\tilde{O}(1/(\Delta_{t_{1}}(a(i,j))-\delta)^{2})=\tilde{O}(1/\delta^{2})$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ and $m^{\mathcal{G}}=\lceil 2V_{T}/\delta\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=(1/C(2\delta)^{\beta})-1$ for some constant $C>0$, we have
@@
Case 2: Now we consider $V_{T}\leq\max\{1/\sqrt{T},1/T^{1/(\beta+1)}\}$ in the following. In this case, we have $\delta=c_{1}\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}.$ For getting $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}$, here we define the policy $\pi$ after time $T$ such that it pulls $V_{T}$ amount of rotting variation for a good arm and 00 for a bad arm. We note that defining how $\pi$ works after $T$ is only for the proof to get a regret bound over time horizon $T$. For the last arm $\tilde{a}$ over the horizon $T$, it pulls the arm up to $V_{T}$ amount of rotting variation if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the amount of rotting variation from pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
Under $E_{1}$, when $m^{\mathcal{G}}=C_{3}$ for some constant $C_{3}>0$, we have R^B(T)R^B_m^G.
@@
From Lemma A.2, under $E_{1}$ we can find that $V_{i}^{\mathcal{G}}\geq\min\{\delta/2,V_{T}\}$ for $i\in[m^{\mathcal{G}}]$. Then if $m^{\mathcal{G}}=C_{3}$ for large enough $C_{3}>0$, then with $\delta=c_{1}\max\{1/T^{1/(\beta+1)},1/\sqrt{T}\}$ and $V_{T}\leq\max\{1/T^{1/(\beta+1)},1/\sqrt{T}\}$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=c_{1}\max\{(1/T)^{1/(\beta+1)},1/T^{1/2}\}$ and $m^{\mathcal{G}}=C_{3}$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=(1/C(2\delta)^{\beta})-1$ for some constant $C>0$, we have
@@
From Lemma A.1, with $\delta=c_{1}\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}$ we have
@@
From $R^{\pi}(T)=R^{\mathcal{G}}(T)+R^{\mathcal{B}}(T)$ and Lemmas A.1, A.5, A.6 with $\delta=c_{1}\max\{1/T^{\frac{1}{\beta+1}},1/\sqrt{T}\}$ we have